A project that costs $3,100 to install will provide annual cash flows of $810 for each of the next 6 years.


a.

Calculate the NPV if the opportunity cost of capital is 11%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


  NPV $   


b. Is this project worth pursuing?
  • Yes

  • No


c.

What is the project's internal rate of return IRR? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


  IRR %
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=810[1-(1.11)^-6]/0.11

=810*4.23053785

=3426.74

NPV=Present value of inflows-Present value of outflows

=3426.74-3100

=$326.74(Approx).

b.Hence since project has positive NPV;project is worth pursuing(Yes).

c.Let irr be x%
At irr,present value of inflows=present value of outflows.

3100=810/1.0x+810/1.0x^2+...........+810/1.0x^6

Hence x=irr=14.59%(Approx).